Please answer the questions all of them
hy Statement of Admitted Assets, Liabilities & Surplus t of Income & Changes in Surplus Report 5,704 436 245521 2306.2 actsivity in 2009 and 2010. In 2011 CPIC severe claims, esulting in net underwriting The encouraging news for 2011: Aer in of moderate osses for a third straight year. ,3112 33151 2242.22142 ratio of 140%, the claim through Cwith a ratio of 107%, the projected EDERAL TAXES INCURRED Net investment gains in 2011 of $523.489 8320 404, 290,180 (131,013) 315,405 SURPLUS, BECING of the net gains. The resut was a modest addtion to surplus of $115,782 S TO SURPLUS (32.554) t stratep. and overall expense system shouild position CPIC port in 2011. We look forward to success in 242348 SURPLUS ENDING 127,6 8543824 2012. David L Lipoey, Charman CO 16
Financial Ratios Exercise #1 Name: Develop the following information using the Chautauqua Patron Insurance Company Annual Report 1. Written Premium 2. Earned Premium 3. Company Loss Ratio 4. Company Expense Ratio using GAP 5. Combined Ratio 6. Operating Ratio
7. Why are the Written Premium and Earned Premium Different? 8. Did the Company make an overall profit or loss? Explain.
Answer #1

1. Written Premium : A written premium describes the total premiums collected on policies issued by an insurance company during a period of time.,

In other words, Written Premium is the gross sum of Insurance premiums received from the customers.

Written Premium is not available in the given financials, it will be reflected in Cash flow statement.

2. Earned Premium : Earned premium as the word suggests is "EARNED" by the insurer, it is a prorated amount of paid-in-advance premiums that has been earned and now belongs to the insurer.

In other words, the earned premium is the part of an insurance premium that paid for a part of time in which the insurance policy was in effect, but has now passed and expired. Since the insurance company covered the risk already during that passed time, it can now consider the premium payments it took from the customers as "EARNED."

Earned Premium for 2011 =$ 6,036,258.

3. Loss Ratio : Loss ratio describes the proportion of Incurred Losses to Earned Premium.

In other words, it is the proportion of Earned Premium that has been used to pay for underwritten policies.

Loss Ratio = (Incurred Losses + Loss exp incurred )/ Earned Premium =(3057509+952731) / 6036258 = 66.43%.

4. Expense Ratio : Expense Ratio = Proportion of all expenses to premium earned.

=(3057509+962371+2445291)/6036258 = 107.11%.

8. The company did make a profit for the year, even though the losses from expenses from policy unaderwriting is more than the premium earned, profit is made thanks to income from investments of $ 523,489. and Financial Charges on calls of $ 101684

Net Profit made after Taxes is $ 290,180